OpenStudy (anonymous):

Find a quadratic model for the set of values (-2, -20), (0, -4), (4, -20)

OpenStudy (amistre64):

we have one root stated, and mirrored ys

OpenStudy (amistre64):

the axis of symmetry, the mirrord ys, has to be the middle of -2 and 4 and when x=0, ah, the y intercept

OpenStudy (anonymous):

a0^2+b0+c=-4 means c = -4

OpenStudy (anonymous):

You can solve this problem by writing 3 equations. The general quadratic equation is \[ax ^{2}+bx+c=0\] You have 3 points given, can you make the 3 equations?

OpenStudy (amistre64):

y = a(x-1)^2 + k 0 = a(-5)^2+k -20 = a(3)^2+k 3 equations with 2 unknowns is one way to see it

OpenStudy (anonymous):

Am not gettin it

OpenStudy (amistre64):

what have you tried?

OpenStudy (amistre64):

bah .... i mixed x,y parts on my first one :) -4 = a(-1)^2+k -20 = a(3)^2+k

OpenStudy (anonymous):

sorrrrrrrrrrrry \[-2x^2-12x-4\]

OpenStudy (anonymous):

Lilmoe is this clear? I think we got you confused.

OpenStudy (anonymous):

Sorry yea am confuse

OpenStudy (amistre64):

there are many ways to approach a solution so its hard to define a specific method withou knowing what your own thoughts about it are

OpenStudy (amistre64):

in general we have 3 points (\(x_1,~y_1\)); (\(x_2,~y_2\)); (\(x_3,~y_3\)) we can setup an equation that can focus on one aspect at a time ... we know that anything subtracted from itself is equal to 0, and that anything times 0 is also 0 so we want to make use of these properties: (\(x_1-x\)); (\(x_2-x\)); (\(x_3-x\)) we can construct an equation with coefficients: \(c_0,~c_1,~c_2,~...,~c_n\) such that all but one unknown coefficient will zero out and leave us with 1 term to find a solution for like this: \[y=c_0+c_1(x_1-x)+c_2(x_1-x)(x_2-x)\] when x=x1, y=y1 and the c1,c2 terms zero out giving us \[y_1=c_0\] knowing c0, we can let x=x2, y=y2 zeroes out the c3 \[y_2=y_1+c_1(x_1-x_2)\] \[\frac{y_2-y_1}{x_1-x_2}=c_1\] c_1 is just the linear slope between the first 2 points ... letting x=x3 and y=y3 lets us solve the final coefficient \[y_3=y_1+\frac{y_2-y_1}{x_1-x_2}(x_1-x_3)+c_2(x_1-x_3)(x_2-x_3)\] \[y_3-y_1-\frac{y_2-y_1}{x_1-x_2}(x_1-x_3)=c_2(x_1-x_3)(x_2-x_3)\] \[\frac{y_3-y_1}{(x_1-x_3)(x_2-x_3)}-\frac{y_2-y_1}{(x_1-x_2)(x_2-x_3)}=c_2\]